A point mass m 167 g hangs from a string of lengthplease answer and explain!
Answer #1

In this case Total kinetic energy get transformed in to total potential energy. i.e.

\frac{1}{2}m.v^{2} = mgh

From the figure

L.Cose h = L- L.Cos(8)

\frac{1}{2}m.v^{2} = mg(L-L\cos \theta )

\frac{1}{2}.v^{2} = g(L-L\cos \theta )

v^{2} = 2.g(L-L\cos \theta )

v = \sqrt{2.g(L-L\cos \theta )} = \sqrt{2\times 9.8(1.79-1.79\times \cos 39.1^{0})}

v = \sqrt{2\times 9.8 \times (1.79-1.39)} = \sqrt{7.84} = 2.80 m / s

So Option C is correct.